# Improving variational counterdiabatic driving with weighted actions and computer algebra

### v1

@@ Introduction
Here, $w_{n}^{(K)}\equiv K^{2}(\epsilon_{n}-E^{(K)})^{2K-2}$ is the state-dependent weight introduced in Sec. III.1, and we define $\chi_{mn}^{(K)}\coloneqq 1/2$ if $w_{m}^{(K)}=w_{n}^{(K)}$ and $\chi_{mn}^{(K)}\coloneqq 1$ otherwise. This decomposition shows that $\mathcal{S}^{(K)}[V]$ is understood as a weighted sum of the partial actions $\mathcal{T}^{(K,n)}[V]$ with the weight $w_{n}^{(K)}$. The partial action $\mathcal{T}^{(K,n)}[V]$ penalizes those deviations $[V-\Phi]_{mn}$ whose weight $w_{m}^{(K)}$ is smaller than $w_{n}^{(K)}$. The partial action depends on $K$, but the dependence is algebraic with a scaling between $O(K^{0})$ and $O(K^{-2})$ (see Appendix A.5). This algebraic dependence is much weaker than the exponential dependence of the weight $w_{n}^{(K)}$ on $K$. Therefore, the $K$-dependence of $\mathcal{S}^{(K)}[V]$ is predominantly captured by the weight $w_{n}^{(K)}$ in the decomposition in Eq. (32a).
Using this decomposition, we first discuss the behavior of $\mathcal{S}^{(K)}[V]$ in the $K\to\infty$ limit. For a sufficiently large $K$, the weight $w_{n}^{(K)}$ is concentrated on the ground state, and thus, the action $\mathcal{S}^{(K)}[V]$ is approximated by $2\hbar^{-2}w_{1}^{(K)}\mathcal{T}^{(K,1)}[V]$. Furthermore, $\mathcal{T}^{(K,1)}[V]$ is approximated by $K^{-2}(\epsilon_{1}-E^{(K)})^{2}\sum_{m:\,m\neq 1}|[V-\Phi]_{1m}|^{2}$ (see Appendix A.5). Combining these two approximations, we observe that the action $\mathcal{S}^{(K)}[V]$ approaches
@@
up to an overall normalization in the $K\to\infty$ limit (we show another derivation of Eq. (33) in Appendix A.5). This limiting action penalizes only $[V-\Phi]_{1n}$ for $n\geq 2$, i.e., the matrix elements of $V-\Phi$ between the ground and excited states.
@@
In fact, this form of penalty is suitable for suppressing the nonadiabatic transitions from the ground state to the excited states. Escape from the ground state is caused only by $[V-\Phi]_{1n}$ for $n=2,\dots,D$, as seen from the Schrödinger equation in Eq. (3). Moreover, the final fidelity, denoted by $|c_{1}(t_{\mathrm{d}})|^{2}$ in the notation of Eq. (3), obeys the quantum speed limit formula [see Eq. (31) of Ref. [53]]:
[…]
@@
The expression inside the square root in Eq. (34) is equal to $\hbar^{2}\mathcal{S}^{(\infty)}[V]/2$. Since $\arccos|c_{1}(t_{\mathrm{d}})|$ is a decreasing function of $|c_{1}(t_{\mathrm{d}})|$ over $0\leq|c_{1}(t_{\mathrm{d}})|\leq 1$, Eq. (34) shows that the final fidelity is lower-bounded by the time integral of the square root of $\mathcal{S}^{(\infty)}[V]$. In this sense, the action $\mathcal{S}^{(\infty)}[V]$ is a natural penalty function to determine driving coefficients that could maximize $|c_{1}(t_{\mathrm{d}})|$, and therefore we call $\mathcal{S}^{(\infty)}[V]$ the ideal action.
@@
Figure 6: Deviation of the driving coefficients from the “ideal” coefficient $\bm{\alpha}^{(\infty)}$, which minimizes the ideal action $\mathcal{S}^{(\infty)}[V(\bm{\alpha})]$. Dashed horizontal lines denote the deviation $\|\bm{\alpha}^{(K)}-\bm{\alpha}^{(\infty)}\|^{2}_{2}$, where $\bm{\alpha}^{(K)}$ is the driving coefficient obtained from the weighted variational method of degree $K$. Colored dots represent the deviation $\|\bm{\alpha}^{(K,n)}-\bm{\alpha}^{(\infty)}\|^{2}_{2}$ plotted against the energy eigenvalue $\epsilon_{n}$, where $\bm{\alpha}^{(K,n)}$ is the minimizer of the partial action $\mathcal{T}^{(K,n)}[V(\bm{\alpha})]$. We fix $N=12$ and $\lambda=0.25$ as an example. (a) Ferromagnetic system with the one-body driving. (b) Spin-glass system with the one-body driving.
@@
Next, we focus on realistic values of $K$ such as $K\leq 5$. For such a small $K$, the weight $w_{n}^{(K)}$ is not peaked enough to single out the $n=1$ partial action $\mathcal{T}^{(K,1)}[V]$ in Eq. (32a), and it is only mildly concentrated in the broad range of lower-energy states. Nevertheless, this mild concentration could still lead to an improvement in CD driving by the following mechanism. Let us hypothesize that $\mathcal{T}^{(K,n)}[V]$ depends smoothly on $n$, and thus $\mathcal{T}^{(K,n)}[V]$ with a small $n$ is closer to $\mathcal{T}^{(K,1)}[V]$ than $\mathcal{T}^{(K,n)}[V]$ with a large $n$. Then, even a mild concentration of $w_{n}^{(K)}$ in small $n$’s can shift $\mathcal{S}^{(K)}[V]$ towards $\mathcal{T}^{(K,1)}[V]$. The $n=1$ partial action $\mathcal{T}^{(K,1)}[V]$ is expected to be close to the ideal action because it contains the same set of terms as the ideal action ($[V-\Phi]_{1n}$ with $n\geq 2$). Therefore, under the above hypothesis, the weighted action is expected to approach the ideal action as $K$ increases, even for small $K$. This scenario is a possible way to understand the improved final fidelity with small $K$.
@@
This scenario is indeed numerically confirmed with our example systems in Fig. 6. We define $\bm{\alpha}^{(\infty)}$ and $\bm{\alpha}^{(K,n)}$ as the minimizers of $\mathcal{S}^{(\infty)}[V(\bm{\alpha})]$ and $\mathcal{T}^{(K,n)}[V(\bm{\alpha})]$, respectively, and plot the deviation $\|\bm{\alpha}^{(K,n)}-\bm{\alpha}^{(\infty)}\|^{2}_{2}$ by varying $K$ and $n$, where $\|\cdot\|_{2}$ is the 2-norm. The figure shows that the deviation does not depend significantly on $K$, but it has a clear trend with $n$. The deviation is small for the ground state and tends to be larger for a larger $n$. This trend implies that $\mathcal{T}^{(K,n)}[V]$ with a small $n$ is closer to the ideal action than $\mathcal{T}^{(K,n)}[V]$ with a large $n$, consistent with the above scenario. Note that the deviation shows an irregular behavior around $\epsilon_{n}\simeq E^{(K)}$ for $K\geq 2$, but this irregularity does not significantly affect the overall action $\mathcal{S}^{(K)}[V]$ because the overall magnitude of $\mathcal{T}^{(K,n)}[V]$ is small for such $n$. This irregular behavior occurs because the weight $w_{n}^{(K)}$ is close to zero for $\epsilon_{n}\simeq E^{(K)}$, and thus $\mathcal{T}^{(K,n)}[V]$ contains only a small number of terms, making its minimization too sensitive to the individual elements $[V-\Phi]_{mn}$.
@@
This scenario is also confirmed by looking at the deviation between the resulting coefficient $\bm{\alpha}^{(K)}$ from the weighted variational method and the ideal coefficient $\bm{\alpha}^{(\infty)}$. We plot the deviation $\|\bm{\alpha}^{(K)}-\bm{\alpha}^{(\infty)}\|^{2}_{2}$ in Fig. 6 (dashed horizontal lines). The plot shows that the driving coefficient $\bm{\alpha}^{(K)}$ approaches the ideal coefficient $\bm{\alpha}^{(\infty)}$ as $K$ increases, confirming the scenario above even for a small $K$.
@@
Figure 7: Response of the driving coefficient $\alpha_{\mu}^{(K)}$ to a perturbation in the longitudinal magnetic field. We fix $N=15$ ($N_{\mathrm{w}}=5$ and $N_{\mathrm{h}}=3$) and $\lambda=0.25$. The index $\mu$ of the driving terms runs from 1 to 15 for the one-body driving [Eq. (28)]. The index runs from $1$ to $37$ for the two-body driving, in which $\mu\leq 15$ are the coefficients on the $Y_{i}$ terms, and $\mu\geq 16$ are on the $(Y_{i}Z_{j}+Z_{i}Y_{j})$ terms [Eq. (29)]. (a) Indices of the spins ($\mu=1,\dots,15$) and the pairs of neighboring spins ($\mu=16,\dots,37$) used in this figure. We give a perturbation $\delta\ln h_{1}=1.02$ at spin 1 (blue). (b) Response function $\mathcal{R}_{\mu}^{(K)}$ for ferromagnetic systems with the one-body driving. Circle shows the median, and the error bar shows the quartile range of 100 instances. (c) Response function for antiferromagnetic systems with the two-body driving. The median and quartile range over 100 instances are shown.
@@
Next, we demonstrate that the driving coefficients from the weighted variational method reflect the nonlocality of the weighted action. We capture the nonlocal dependence of the driving coefficients on the system parameters by giving a small perturbation to the system parameter and observing the resulting change in the driving coefficients. Concretely, we modify the longitudinal field at site 1 by $\ln h_{1}\to\ln h_{1}+\delta\ln h_{1}$ and compute the resulting change in the driving coefficients $\ln|\alpha_{\mu}^{(K)}|\to\ln|\alpha_{\mu}^{(K)}|+\delta\ln|\alpha_{\mu}^{(K)}|$ for every index $\mu$ of the driving terms. We define the response function as
@@
which varies from instance to instance. Here, we focus on the logarithmic change (fold change) so that the response function is independent of the magnitudes of $h_{1}$ and $|\alpha_{\mu}^{(K)}|$. Taking $N=15$ as an example, we assign the indices $\mu$ as in Fig. 7(a) so that a larger index corresponds to a driving at a farther position from the perturbed spin.
@@
For the two-body driving, the response functions are nonzero for any $\mu$ and any $K\geq 1$, but their magnitudes differ significantly between $K=1$ and $K\geq 2$ [Fig. 7(c)]. The response functions for $K=1$ drop exponentially with the spatial distance between the perturbed spin and the site on which the driving coefficient acts. On the other hand, the response functions from the weighted variational method ($K\geq 2$) remain at significantly higher values even when the distance increases. This difference implies that the weighted variational method can incorporate nonlocal information much more efficiently than the conventional method.
@@
In this paper, we have proposed the weighted variational method to derive refined variational AGPs for general parameterized Hamiltonians, which significantly improves variational CD driving. The proposal combines a theoretical advance and a numerical algorithm (Sec. II). On the theoretical side, we find that the exact AGP has an infinite variety of algebraic characterizations involving fictitious Hamiltonians $\mathcal{P}_{\lambda}(H(\lambda))$, which produces an infinite number of weighted actions. Minimizing a weighted action under an operator ansatz gives an improved CD driving protocol. We have also developed a computer algebra algorithm to perform the minimization efficiently in polynomial time with respect to the system size.
@@
We have applied our proposal to ground-state evolution (Sec. III) and demonstrated that it significantly improves the final fidelity in quantum Ising models (Sec. IV). We have analyzed its possible mechanisms from two aspects in Sec. V. First, the weighted action emphasizes the partial actions associated with lower-energy states, whose minimizers are closer to the minimizer of the ideal action. The ideal action is a suitable penalty function for suppressing nonadiabatic transitions from the ground state, as understood by the quantum speed limit formula. Second, due to the nonlocality of the weighted action, the driving coefficients depend on the system parameters in a nonlocal way.
@@
Our proposal applies to a wide range of systems, and it can update almost all previous uses of conventional variational CD driving by replacing the conventional action with a weighted action. In particular, our concrete C++ implementation of the ground state evolution of spin-$1/2$ systems can be readily applied to various practical problems. It can be used to improve the preparation of nontrivial quantum states, such as the GHZ state [57] and the Kitaev ground state [32], as well as to enhance computational tasks such as optimization [43, 44, 45, 46, 47, 48] and adiabatic quantum computation. Our theoretical framework should also be straightforwardly applicable to the ground-state evolution of fermionic systems after implementing a computer algebra for fermions. A more nontrivial future application is to quantum processes other than ground-state evolution. Such processes include the adiabatic preparation of the highest energy state and a mid-spectrum eigenstate [36], as well as the adiabatic evolution of a Gibbs state to assist heat engines [38, 39]. These processes would be improved by designing weighted actions that emphasize the relevant parts of the energy spectrum. For example, a weighted action for preparing a mid-spectrum eigenstate could be designed by regarding the problem as the low-energy state preparation of $[H(\lambda)-e_{\lambda}]^{2}$, where a constant $e_{\lambda}$ is chosen close to the energy of the target state, and using our framework of ground-state evolution in Sec. III.
@@
The weighted variational method can be combined with various existing frameworks to further improve CD driving. For example, some literature considers optimizations of the intermediate driving path between fixed initial and final states by equipping the Hamiltonian with extra degrees of freedom to be optimized [31, 56, 57]. This approach can be easily combined with the weighted variational method. Although the extra optimization requires the computation of driving coefficients multiple times, this computational cost can be reduced by separating the parameter-independent traces of operators with parameter-dependent coefficients, as in Sec. II.3. It is also an interesting direction to incorporate weighted actions into existing frameworks with tensor networks [51, 52], digital quantum simulation [43, 44], and Floquet-based methods [49, 50].
@@
From a computational point of view, we leave two problems for future work. First, as mentioned above, a computer algebra framework for fermionic systems should be explicitly constructed and implemented. This can be achieved by combining an existing treatment of fermionic operators based on normal-order products [34] with our algorithm for spin-$1/2$ systems. Second, while our algorithm works for systems without any regular structure, it may be possible to find a more efficient algorithm by taking advantage of regular lattice structures. In fact, the traces of the powers of the Hamiltonian $\mathop{\mathrm{tr}}(H^{k})$ for $k=1,2,\dots$ for such systems can be efficiently evaluated using graph-theoretic considerations, as used in the high-temperature expansion in statistical mechanics [77, 78]. Similar considerations may be useful for evaluating $Q^{(\mathcal{P})}_{\mu\nu}(\lambda)$ and $r^{(\mathcal{P})}_{\mu}(\lambda)$ more efficiently.
@@
While we have focused entirely on the application of variational AGPs to CD driving, our results may also have implications for the role of variational AGPs as probes of fundamental properties of quantum many-body systems. The exact AGP, along with closely related concepts such as the quantum geometric tensor [21, 15] and the regularized AGP [79], is known to be a sensitive probe of quantum phase transitions [80] and quantum chaos [79]. Inheriting this role, variational AGPs have been used to probe phase transitions [53, 55], quantum chaos [54], and macroscopic singularities [67]. These roles of variational AGPs could be enhanced by the weighted variational method. For example, the variational AGP obtained from the weighted variational method for ground-state evolution (Sec. III) should reflect the properties of the ground state more sensitively than the conventional method, and thus it may be more sensitive to quantum phase transitions, which are most prominent in the ground state. Exploring the impact of the weighted variational method in these fundamental contexts is another interesting direction for future research.

### v3

@@ Introduction
Here, $w_{n}^{(K)}\equiv K^{2}(\epsilon_{n}-E^{(K)})^{2K-2}$ is the state-dependent weight introduced in Sec. III.1, and we define $\chi_{mn}^{(K)}\coloneqq 1/2$ if $w_{m}^{(K)}=w_{n}^{(K)}$ and $\chi_{mn}^{(K)}\coloneqq 1$ otherwise. This decomposition shows that $\mathcal{S}^{(K)}[V]$ is understood as a weighted sum of the partial actions $\mathcal{T}^{(K,n)}[V]$ with the weight $w_{n}^{(K)}$. The partial action $\mathcal{T}^{(K,n)}[V]$ penalizes those deviations $V-\Phi]_{mn}$ whose weight $w_{m}^{(K)}$ is smaller than $w_{n}^{(K)}$. The partial action depends on $K$, but the dependence is merely algebraic, with a scaling between $O(K^{0})$ and $O(K^{-2})$ (see Appendix B.4). This algebraic dependence is much weaker than the exponential dependence of the weight $w_{n}^{(K)}$ on $K$. Therefore, the $K$-dependence of $\mathcal{S}^{(K)}[V]$ is predominantly captured by the weight $w_{n}^{(K)}$ in the decomposition in Eq. (31a).
Using this decomposition, we first discuss the behavior of $\mathcal{S}^{(K)}[V]$ in the $K\to\infty$ limit. For a sufficiently large $K$, the weight $w_{n}^{(K)}$ is concentrated on the ground state, and thus, the action $\mathcal{S}^{(K)}[V]$ is approximated by $2\hbar^{-2}w_{1}^{(K)}\mathcal{T}^{(K,1)}[V]$. Furthermore, $\mathcal{T}^{(K,1)}[V]$ is approximated by $K^{-2}(\epsilon_{1}-E^{(K)})^{2}\sum_{m:\,m\neq 1}|[V-\Phi]_{1m}|^{2}$ (see Appendix B.4). Combining these two approximations, we observe that the action $\mathcal{S}^{(K)}[V]$ approaches
@@
up to an overall normalization and a $V$-independent constant term in the $K\to\infty$ limit (we show another derivation of Eq. (32) in Appendix B.4). This limiting action penalizes only $V-\Phi]_{1n}$ for $n\geq 2$, i.e., the matrix elements of $(V-\Phi)$ between the ground and excited states.
@@
In fact, this penalty function in Eq. (32) is suitable for suppressing nonadiabatic transitions from the ground state to the excited states. The escape of the population from the ground state is caused only by $V-\Phi]_{1n}$ with $n\geq 2$, as seen from the Schrödinger equation in Eq. (3). Moreover, the final fidelity, denoted by $|c_{1}(t_{\mathrm{d}})|^{2}$ in the notation of Eq. (3), obeys the quantum speed limit formula [see Eq. (31) of Ref. [57]]:
$\arccos|c_{1}(t_{\mathrm{d}})|\leq\int_{0}^{t_{\mathrm{d}}}\frac{|\dot{\lambda}_{t}|}{\hbar}\sqrt{\sum_{n\neq 1}\bigl|[V(\lambda_{t})-\Phi(\lambda_{t})]_{1n}\bigr|^{2}}dt.$ (33)
@@
The expression inside the square root in Eq. (33) is equal to $\hbar^{2}\mathcal{S}^{(\infty)}[V]/2$. Since $\arccos x$ is a decreasing function over $0\leq x\leq 1$, Eq. (33) provides a lower bound of the final fidelity $|c_{1}(t_{\mathrm{d}})|$ in terms of $S^{(\infty)}[V]$. It shows that if the time integral of $\sqrt{\mathcal{S}^{(\infty)}[V]}$ is smaller, the final fidelity is closer to unity. In this sense, the action $\mathcal{S}^{(\infty)}[V]$ is a natural penalty function to determine driving coefficients that could maximize $|c_{1}(t_{\mathrm{d}})|$. We call $\mathcal{S}^{(\infty)}[V]$ the ideal action for convenience.
@@
Figure 7: Deviation of the driving coefficients from the “ideal” coefficient $\bm{\alpha}^{(\infty)}$, which minimizes the ideal action $\mathcal{S}^{(\infty)}[V(\bm{\alpha})]$. Colored dots represent the deviation $\|\bm{\alpha}^{(K,n)}-\bm{\alpha}^{(\infty)}\|^{2}_{2}$ plotted against the energy eigenvalue $\epsilon_{n}$, where $\bm{\alpha}^{(K,n)}$ is the minimizer of the partial action $\mathcal{T}^{(K,n)}[V(\bm{\alpha})]$. Dashed horizontal lines denote the deviation $\|\bm{\alpha}^{(K)}-\bm{\alpha}^{(\infty)}\|^{2}_{2}$, where $\bm{\alpha}^{(K)}$ is the driving coefficient obtained from the weighted variational method of degree $K$. We fix $N=12$ and $\lambda=0.25$ as an example. (a) Ferromagnetic system with the one-body driving. (b) Spin-glass system with the one-body driving.
@@
Next, we focus on realistic values of $K$ such as $K\leq 5$. For such a small $K$, the weight $w_{n}^{(K)}$ is not peaked enough to single out the $n=1$ partial action $\mathcal{T}^{(K,1)}[V]$ in Eq. (31a), and it is only mildly concentrated in the broad range of lower-energy states. Nevertheless, this mild concentration could still lead to an improvement in CD driving by the following mechanism. Let us hypothesize that $\mathcal{T}^{(K,n)}[V]$ depends smoothly on $n$, and thus $\mathcal{T}^{(K,n)}[V]$ with small $n$ is closer to $\mathcal{T}^{(K,1)}[V]$ than $\mathcal{T}^{(K,n)}[V]$ with large $n$. Then, even a mild concentration of $w_{n}^{(K)}$ in small $n$’s can shift $\mathcal{S}^{(K)}[V]$ towards $\mathcal{T}^{(K,1)}[V]$. The $n=1$ partial action $\mathcal{T}^{(K,1)}[V]$ is expected to be close to the ideal action because it contains the same set of terms as the ideal action ($V-\Phi]_{1n}$ with $n\geq 2$). Therefore, under the above hypothesis, the weighted action is expected to approach the ideal action as $K$ increases, even for small $K$. This scenario provides a possible way to understand the improved final fidelity with small $K$.
@@
This scenario is indeed numerically confirmed with our example systems (Fig. 7). We define $\bm{\alpha}^{(\infty)}$ and $\bm{\alpha}^{(K,n)}$ as the minimizers of $\mathcal{S}^{(\infty)}[V(\bm{\alpha})]$ and $\mathcal{T}^{(K,n)}[V(\bm{\alpha})]$, respectively, and plot the deviation $\|\bm{\alpha}^{(K,n)}-\bm{\alpha}^{(\infty)}\|^{2}_{2}$ by varying $K$ and $n$, where $\|\cdot\|_{2}$ is the 2-norm. The figure shows that the deviation does not depend significantly on $K$, but it has a clear trend with $n$. The deviation is small for the ground state and tends to be larger for a larger $n$. This trend implies that $\mathcal{T}^{(K,n)}[V]$ with small $n$ is closer to the ideal action than $\mathcal{T}^{(K,n)}[V]$ with large $n$, consistent with the above scenario. Note that the deviation shows an irregular behavior around $\epsilon_{n}\simeq E^{(K)}$ for $K\geq 2$, but this irregularity does not significantly affect the overall action $\mathcal{S}^{(K)}[V]$ because the overall magnitude of $\mathcal{T}^{(K,n)}[V]$ is small for such $n$. This irregular behavior occurs because the weight $w_{n}^{(K)}$ is close to zero when $\epsilon_{n}\simeq E^{(K)}$, and thus $\mathcal{T}^{(K,n)}[V]$ contains only a small number of terms, making its minimization too sensitive to individual elements $V-\Phi]_{mn}$.
@@
The scenario is also confirmed by looking at the deviation between the coefficient $\bm{\alpha}^{(K)}$ from the weighted variational method, which is the minimizer of $\mathcal{S}^{(K)}[V(\bm{\alpha})]$, and the ideal coefficient $\bm{\alpha}^{(\infty)}$. We plot the deviation $\|\bm{\alpha}^{(K)}-\bm{\alpha}^{(\infty)}\|^{2}_{2}$ in Fig. 7 (dashed horizontal lines). The plot shows that the driving coefficient $\bm{\alpha}^{(K)}$ approaches the ideal coefficient $\bm{\alpha}^{(\infty)}$ as $K$ increases, confirming the scenario above even for a small $K$.
@@
Figure 8: Response of the driving coefficient $\alpha_{\mu}^{(K)}$ to a perturbation in the longitudinal magnetic field. We fix $N=15$ ($N_{\mathrm{w}}=5$ and $N_{\mathrm{h}}=3$) and $\lambda=0.25$. The index $\mu$ of the driving terms runs from 1 to 15 for the one-body driving [Eq. (27)]. The index runs from $1$ to $37$ for the two-body driving, in which $\mu\leq 15$ are the coefficients on the $Y_{i}$ terms, and $\mu\geq 16$ are on the $(Y_{i}Z_{j}+Z_{i}Y_{j})$ terms [Eq. (28)]. (a) Indices of the spins ($\mu=1,\dots,15$) and the pairs of neighboring spins ($\mu=16,\dots,37$) used in this figure. We give a perturbation $\delta\ln h_{1}=\ln 1.02$ at spin 1 (blue). (b) Absolute value of the response function $|\mathcal{R}_{\mu}^{(K)}|$ for ferromagnetic systems with the one-body driving. Circle shows the median, and the error bar shows the quartile range of 100 instances. (c) Absolute value of the response function for antiferromagnetic systems with the two-body driving. The median and quartile range over 100 instances are shown.
@@
Next, we demonstrate that the driving coefficients from the weighted variational method reflect the nonlocality of the weighted action. We capture the nonlocal dependence of the driving coefficients on the system parameters by giving a small perturbation to a system parameter and observing the resulting change in the driving coefficients. Concretely, we modify the longitudinal field at site 1 by $\ln h_{1}\to\ln h_{1}+\delta\ln h_{1}$ and compute the resulting change in the driving coefficients $\ln|\alpha_{\mu}^{(K)}|\to\ln|\alpha_{\mu}^{(K)}|+\delta\ln|\alpha_{\mu}^{(K)}|$ for every index $\mu$ of the driving terms. We define the response function as
@@
which varies from instance to instance. Here, we focus on the logarithmic change (fold change) so that the response function is independent of the magnitudes of $h_{1}$ and $|\alpha_{\mu}^{(K)}|$. Taking $N=15$ as an example, we assign the indices $\mu$ as in Fig. 8(a) so that a larger index corresponds to a driving field at a farther position from the perturbed spin.
@@
For the two-body driving, the response functions are nonzero for any $\mu$ and any $K\geq 1$, but their magnitudes differ significantly between $K=1$ and $K\geq 2$ [Fig. 8(c)]. The response functions for $K=1$ drop exponentially with the spatial distance between the perturbed spin and the site on which the driving field acts. On the other hand, the response functions from the weighted variational method ($K\geq 2$) remain at significantly higher values even when the distance increases. This difference implies that the weighted variational method can incorporate nonlocal information much more efficiently than the conventional method.
@@
In this paper, we have proposed the weighted variational method to derive refined variational AGPs for general parameterized Hamiltonians, which significantly improves variational CD driving. The proposal combines a theoretical advance and a numerical algorithm (Sec. II). On the theoretical side, we find that the exact AGP has an infinite variety of algebraic characterizations involving fictitious Hamiltonians $\mathcal{P}_{\lambda}(H(\lambda))$, which results in an infinite number of weighted actions. Minimizing a weighted action under an operator ansatz gives an improved CD driving protocol. We have also developed a computer algebra algorithm to perform the minimization efficiently in polynomial time with respect to the system size.
@@
We have applied our proposal to ground-state evolution (Sec. III) and demonstrated that it significantly improves the final fidelity in quantum Ising models (Sec. IV). The median improvement in final fidelity is more than tenfold for $N=15$ for all three classes of systems, and this improvement is robust across the vast majority of randomly generated instances, except for five out of 300 spin-glass instances. We have analyzed its possible mechanisms from two aspects in Sec. V. First, the weighted action emphasizes the partial actions associated with lower-energy states, whose minimizers are closer to the minimizer of the ideal action. The ideal action is a suitable penalty function for suppressing nonadiabatic transitions from the ground state, as understood by the quantum speed limit formula. Second, due to the nonlocality of the weighted action, the driving coefficients depend on the system parameters in a nonlocal way.
@@
Our proposal is applicable to a wide range of systems, and we expect that it can improve almost all previous uses of conventional variational CD driving. In particular, our concrete C++ implementation of the ground state evolution of spin-$1/2$ systems can be readily applied to various practical problems. It can be used to improve the preparation of nontrivial quantum states, such as the GHZ state [61] and the Kitaev ground state [32], as well as to enhance computational tasks such as optimization [44, 45, 46, 47, 48, 49] and adiabatic quantum computation. Testing these applications with a greater number of spins than those used in Sec. IV is an important future work. Our theoretical framework should also be straightforwardly applicable to the ground-state evolution of fermionic systems after implementing a computer algebra for fermions. A more nontrivial future application is to quantum processes other than ground-state evolution. Such processes include the adiabatic preparation of the highest energy state and a mid-spectrum eigenstate [37], as well as the adiabatic evolution of a Gibbs state to assist heat engines [39, 40]. These processes would be improved by designing weighted actions that emphasize the relevant parts of the energy spectrum. For example, a weighted action for preparing a mid-spectrum eigenstate could be designed by regarding the problem as the preparation of the low-energy state of $H(\lambda)-e_{\lambda}]^{2}$, where a constant $e_{\lambda}$ is chosen close to the energy of the target state, and using our framework of ground-state evolution in Sec. III.
@@
The weighted variational method can be combined with various existing frameworks to further improve CD driving. For example, some literature considers optimizations of the intermediate driving path between fixed initial and final states by equipping the Hamiltonian with extra degrees of freedom to be optimized [31, 60, 61]. This approach can be easily combined with the weighted variational method. Although the extra optimization requires us to compute driving coefficients multiple times, this computational cost can be reduced by separating parameter-independent traces of operators with parameter-dependent coefficients, as has been discussed in Sec. II.4. It is also an interesting direction to incorporate weighted actions into existing frameworks with tensor networks [55, 56], digital quantum simulation [44, 45], and Floquet-based methods [50, 51, 52, 53, 54]. In particular, the Floquet engineering can produce effective coupling terms that are hard to implement directly. The coefficients on these effective terms can be improved using our method.
@@
From a computational point of view, we leave two problems for future work. First, as mentioned above, a computer algebra framework for fermionic systems should be explicitly constructed and implemented. This can be achieved by combining an existing treatment of fermionic operators based on normal-order products [35] with our algorithm for spin-$1/2$ systems. Second, while our algorithm works for systems without any regular structure, it may be possible to find a more efficient algorithm by taking advantage of regular lattice structures. In fact, the traces of the powers of the Hamiltonian $\mathop{\mathrm{tr}}(H^{k})$ for $k=1,2,\dots$ for such systems have been efficiently evaluated using graph-theoretic considerations in high-temperature expansion of Gibbs states [85, 86]. Similar considerations may be useful for evaluating $Q^{(\mathcal{P})}_{\mu\nu}(\lambda)$ and $r^{(\mathcal{P})}_{\mu}(\lambda)$ more efficiently.
@@
While we have focused entirely on the application of variational AGPs to CD driving, our results may also have implications for the roles of variational AGPs as probes of fundamental properties of quantum many-body systems. The exact AGP, along with closely related concepts such as the quantum geometric tensor [21, 15] and the regularized AGP [87], is known to be a sensitive probe of quantum phase transitions [88] and quantum chaos [87]. Inheriting this role, variational AGPs have been used to probe phase transitions [57, 59], quantum chaos [58], and macroscopic singularities [71]. These roles of variational AGPs could be enhanced by the weighted variational method. For example, the variational AGP obtained from the weighted variational method for ground-state evolution (Sec. III) should reflect the properties of the ground state more sensitively than the conventional method, and thus it may be more sensitive to quantum phase transitions, which are most prominent in the ground state. Exploring the impact of the weighted variational method in these fundamental contexts is another interesting direction for future research.